# Chinese Academy of Sciences, Beijing 100190, China

### v1

@@ 4.2 Results of BBN abundances
In this subsection we present the resulting light element abundances incorporating all the corrections discussed in the last subsection. To do this, we have added the aforementioned correction induced by the neutrino spectral distortions in the public BBN code PRIMATPitrou:2018cgg and solved the abundances of the light elements. For the computation we have used the latest values of other physical and cosmological parameters, $e.g.,$ the neutron lifetime is taken as $\tau_{n}=879.4s$, the axial coupling of the nucleons is $g_{A}=1.2756$, and the baryon density is $\omega_{b}h^{2}=0.024$. For convenience, we define the mass fraction of isotope $i$ as $X_{i}\equiv A_{i}\frac{n_{i}}{n_{b}}$, where $A_{i}$ and $n_{i}$ are the baryon number and the number density of isotope $i$ and $n_{b}$ is the total number density of baryon. Customarily, we also denote $Y_{P}\equiv X_{{}^{4}{\rm He}}$ and $i/\rm H\equiv n_{i}/n_{{}^{1}\rm H}$.
@@
Figure 5: Light element abundances as a function of the neutrino degeneracy parameter $\xi_{\nu}$, normalized to the results for $\xi_{\nu}=0$. The ${}^{3}{\rm He}$ stands for ${}^{3}{\rm He}+\rm T$ and ${}^{7}{\rm Li}$ stands for ${}^{7}{\rm Li}+\rm^{7}Be$ due to the slow radioactive decays. All the optional weak rate corrections in PRIMAT (see Ref. Pitrou:2018cgg ) and the new correction from neutrino spectral distortions are included in the calculation. The $68\%$ constraints on the helium abundance $Y\rm_{P}$ from the CMB (${Y\rm_{P}}=0.242\pm 0.012$) and the recent EMPRESS experiment (${Y\rm_{P}}=0.2370^{+0.0034}_{-0.0033}$) are indicated by the light and dark blue shaded regions, respectively.
@@
We show in Fig. 5 the resulting light element abundances, including all the aforementioned weak rate corrections in (79) and the QED corrections up to $\mathcal{O}(e^{3})$ for plasma thermodynamics and certain nuclear rates, as a function of the neutrino degeneracy parameter $\xi_{\nu}$. The abundance of ${}^{3}{\rm He}$ stands for ${}^{3}{\rm He}+\rm T$ and ${}^{7}{\rm Li}$ for ${}^{7}{\rm Li}+\rm^{7}Be$ to account for the slow radioactive decays. Since the abundances for different elements can vary by several orders of magnitude, we normalize the element abundances by their corresponding abundance with $\xi_{\nu}=0$. Therefore all curves in Fig. 5 intersect at $\xi_{\nu}=0$. We see that the abundances of both $\rm D,\,^{3}{\rm He},\,^{4}{\rm He}$ and ${}^{7}{\rm Li}$ decrease with increasing $\xi_{\nu}$ due to the fact that the larger the $\xi_{\nu}$, the smaller the number density ratio between neutron and proton. In addition, the variation of the ${}^{4}{\rm He}$ abundance is the most sensitive to the variation of $\xi_{\nu}$, while the ${}^{3}{\rm He}$ abundance is the least sensitive. We also show in Fig. 5 the constraints on the helium abundance from obtained by CMB and EMPRESS experiments,
${Y\rm_{P}}|_{\rm CMB}=0.242\pm 0.012,\,\,{Y\rm_{P}}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033},$ (82)
@@
by the light and dark blue shaded regions. From our results, the constraints on ${Y\rm_{P}}$ within the $68\%$ confidence interval can be easily translated into constraints on the neutrino degeneracy parameter $\xi_{\nu}$ as
${\rm CMB}:\,\,-0.026\leq$ $\,\xi_{\nu}\leq 0.069\,,$ (83)
EMPRESS:  ${\rm EMPRESS}:\,\,0.027\leq$ $\,\xi_{\nu}\leq 0.054\,.$ (84)
@@
Clearly, the EMPRESS experiment prefer a nonzero positive $\xi_{\nu}$ while the CMB experiment is still compatible with the SM case $\xi_{\nu}=0$.
@@
Figure 6: The relative difference of the light element abundances compared to the results in the neutrino instantaneous decoupling limit as a function of the neutrino degeneracy parameter $\xi_{\nu}$. The solid and dotted curves correspond to the results with and without the weak corrections induced by the neutrino spectral distortions, respectively. The meaning of the color for the curves is the same as in the plot Fig. 5, except that the negative results of ${}^{7}{\rm Li}$ are shown here for simplicity.
@@
Table 1: Light element abundances including all weak rate corrections in (79), for various implementations of neutrino decoupling and degeneracy parameters. ${}^{3}{\rm He}$ stands for $(^{3}{\rm He}+\mathrm{T})$ and ${}^{7}{\rm Li}$ stands for $(^{7}{\rm Li}+\rm Be)$ to account for slow radioactive decays.
@@
Table 2: Relative difference of the light element abundances for the case with and without the corrections induced by the neutrino spectral distortions compared to the results in the neutrino instantaneous decoupling limit. ${}^{3}{\rm He}$ stands for $(^{3}{\rm He}+\mathrm{T})$ and ${}^{7}{\rm Li}$ stands for $(^{7}{\rm Li}+\rm Be)$ to account for slow radioactive decays.
@@
It is also interesting to discuss the effects of the corrections induced by the neutrino spectral distortions on BBN. Focusing on the corrections induced by different treatments of the neutrino decoupling process, we consider the following three implementations: (1) The neutrino instantaneous decoupling limit, $i.e.,$ the $N_{\rm eff}^{\rm inst}$ as discussed in the last section and the thermal Fermi-Dirac distributions for neutrinos and antineutrinos are used in the calculation of the BBN; (2) The no-SD case, where the actual $N_{\rm eff}$ and neutrino temperatures are used, including the non-instantaneous decoupling effect, but considering only the thermal Fermi-Dirac distributions for neutrinos and antineutrinos. This is the most common treatment for BBN in the studies to date; (3) The actual SD case, using the actual $N_{\rm eff}$ with the non-instantaneous decoupling effect and the actual density distributions with the spectral distortions for neutrinos and antineutrinos, corresponding to the results shown in Fig. 5. Using the BBN results in the instantaneous decoupling limit as the benchmark, we show the relative differences of the results for the no-SD and the actual-SD cases with the dotted and solid curves in Fig. 6, respectively. Note that we have plotted the opposite of the relative differences for ${}^{7}{\rm Li}$ in Fig. 6 as it is negative due to the particular nuclear reaction rates used in PRIMAT. We also report the explicit values of the abundances of helium-4, deuterium, helium-3 and lithium-7 for the degeneracy parameters $\xi_{\nu}=0,\pm 1$ in table 1, and the associated relative variations compared to the baseline abundance in table 2. We also note that our results with $\xi_{\nu}=0$ are consistent with the results reported in Froustey:2020mcq , where they also considered the effects of neutrino spectral distortions to BBN within the SM case.
As we can see in Fig. 6 and the tables, the actual-SD case and the no-SD case yield almost the same light element abundances for $\xi_{\nu}=-0.1$. And the variation of the BBN abundances from the results in the instantaneous decoupling limit for the actual-SD case show little dependence on $\xi_{\nu}$, while the traditional treatment without the spectral distortions predicts a stronger dependence for both elements. In particular, the no-SD case predicts negative variations for the helium abundance with positive $\xi_{\nu}$ compared to the result in the instantaneous decoupling limit. However, the more precise results with the spectral distortions show that the variations of the helium abundance are positive, showing that the inclusion of the neutrino spectral distortions is necessary for a precise and accurate prediction of the BBN abundances.
@@
Finally, we also provide a set of fitting functions for the light element abundances as
@@
where the subscript $\xi_{\nu}=0$ labels the results obtained in the SM case $\xi_{\nu}=0$ using the actual density matrices, see the explicit values in Table 1. Such fitting functions benefit for further studies, and the fractional errors are smaller than $0.1\%$ with the range $-0.1\leq\xi_{\nu}\leq 0.1$.  [^3] The coefficients of our fitting functions are slightly different from the results in Pitrou:2018cgg ; Froustey:2024mgf , which we think is due to the different corrections to the weak interaction rates we considered.
5 Impacts of the Primordial Neutrino Asymmetry on CMB and LSS
In this section we focus on the impacts of primordial neutrino asymmetries on the evolution of the late universe, especially the cosmic microwave background (CMB) and the large scale structure (LSS). In general, the primordial neutrino asymmetry influence CMB and LSS in three ways.
@@ 5 Impacts of the Primordial Neutrino Asymmetry on CMB and LSS
    The expansion rate of the Universe is changed by the increased $N_{\rm eff}$ due to the primordial neutrino asymmetries. Particularly, the time of matter-radiation equality is delayed by the additional radiation density, leading to an enhanced early integrated Sachs-Wolfe effect for CMB angular power spectrum.
@@
    As we have seen in Fig. 5, the abundance of helium is significantly modified by the electron neutrino asymmetry, which will in turn alter the evolution of the late Universe, such as the tail of the CMB angular power spectrum by the amount of diffusion damping and the baryon acoustic oscillation (BAO).
@@
    In addition, in order to incorporate the effect of neutrino oscillation in the process of neutrino decoupling, the mass of neutrinos must be considered. The neutrino masses will alter the evolution of the late universe by such as, affect the matter-radiation equality, even in the SM case with the thermal Fermi-Dirac distributions. For the SM case, the experiment on CMB and BAO can put a relative stringent bound on the sum of neutrino masses $\sum m_{\nu}<0.07\,{\rm eV}$ DESI:2024mwx , we will return to the constraint on neutrino masses in the next section. Here we emphasize that the neutrino Boltzmann hierarchy for massive neutrinos will be furthermore modified by the neutrino degeneracy parameter and the spectral distortions.
In the following we first review the impact of massive neutrinos on the universe and present the Boltzmann hierarchies of electrons, photons and neutrinos. Then we discuss the effects of the primordial neutrino asymmetries on the CMB temperature anisotropies and the linear matter power spectrum computed with the Einstein-Boltzmann code CLASS Lesgourgues:2011re ; Lesgourgues:2011rh .
5.1 Effects of the primordial neutrino asymmetries in structure formation
To describe the flavor oscillations in the neutrino decoupling process, the mass of the neutrinos must be taken into account, which in turn has significant implications for the process of structure formation, since massive neutrinos are the only species in the SM that scale as radiation in the early universe and as matter in the late universe. In the early universe, neutrinos and antineutrinos remain relativistic after decoupling as long as their effective temperatures are much larger than their masses. In this case, the neutrino and antineutrino energy densities contribute to the total radiation energy density, parameterised by the aforementioned $N_{\rm eff}$. Since the neutrino effective temperatures decrease as $a^{-1}$ with the expansion of the universe, neutrinos and antineutrinos become non-relativistic when their effective temperatures fall below their masses, and contribute to matter after that. The non-relativistic transitions and the free-steaming effects of massive neutrinos then leave significant imprints on the CMB anisotropies and the LSS. As mentioned above, the situation is modified in the presence of the primordial neutrino asymmetries. In addition to the indirect change induced by the modified Hubble rate and the helium abundance, we highlight several other modifications due to the primordial neutrino asymmetries in the following.
@@ 5.1.1 Effects on neutrino abundance
First, the total relic abundance of neutrinos and antineutrinos are directly increased due to the primordial neutrino asymmetries. At present, the total neutrino relic abundance can be written as
@@
where $y=p/T_{\rm CM}$ is the comoving momentum defined in section 3, $T_{\rm CM,0}$ is the current comoving temperature, $\rho_{\rm crit}$ is the critical energy density, $\varrho_{i}(y)$ and $\bar{\varrho}_{i}(y)$ are the respective comoving density distributions for neutrinos and antineutrinos with $i=1,2,3$, corresponding to the three mass eigenstates with mass $m_{i}$. Obviously, the present neutrino abundance is also modified by the spectral distortions of neutrinos and antineutrinos, and the integrals over the neutrino distributions cannot be carried out analytically in this case. The effect of the primordial neutrino asymmetries will be clearer by neglecting the spectral distortions and performing the integral analytically, as seen in the second row of (89). The first factor in the second row $\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}$ is the usual neutrino relic abundance in the SM case, while the second factor represents the effect of neutrino asymmetries with ${\rm Li}_{j}(x)$ the polylogarithms of order $j$ and $\zeta(x)$ the Riemann zeta function. For the SM case one finds ${\rm Li}_{3}(-1)=3\zeta(3)/4$, so the second factor becomes 1. For non-zero $\xi_{\nu}$, the second factor depends only on the absolute value of $\xi_{\nu}$, and the larger the $\xi_{\nu}$, the larger the second factor and the larger the neutrino abundance. However, we note that such corrections from primordial neutrino asymmetries are quite small. For example, the resulting relative difference of $\Omega_{\nu}h^{2}$ for $\xi_{\nu}=0.2$ is only $1.54\%$. Similarly, the corrections to the neutrino abundance induced by the spectral distortions of neutrinos and antineutrinos are also negligible.

### v2

@@ 4.2 Results of BBN Abundances
In this subsection, we present the resulting light element abundances, incorporating all corrections discussed previously. Specifically, we have implemented the corrections induced by neutrino spectral distortions into the public BBN code PRIMAT Pitrou:2018cgg and computed the abundances of the light elements. For this computation, we utilized the latest values of relevant physical and cosmological parameters; for instance, the neutron lifetime is taken to be $\tau_{n}=879.4\,\text{s}$, the axial coupling of the nucleons is $g_{A}=1.2756$, and the baryon density is $\omega_{b}h^{2}=0.024$. For convenience, we define the mass fraction of isotope $i$ as $X_{i}\equiv A_{i}\frac{n_{i}}{n_{b}}$, where $A_{i}$ and $n_{i}$ are the baryon number and number density of isotope $i$, respectively, and $n_{b}$ is the total baryon number density. We also customarily denote $Y_{P}\equiv X_{{}^{4}{\rm He}}$ and $i/\rm H\equiv n_{i}/n_{{}^{1}\rm H}$.
@@
Figure 5: Light element abundances as a function of the neutrino degeneracy parameter $\xi_{\nu}$, normalized to the results for $\xi_{\nu}=0$. The ${}^{3}{\rm He}$ notation refers to ${}^{3}{\rm He}+\rm T$, and ${}^{7}{\rm Li}$ denotes ${}^{7}{\rm Li}+\rm^{7}Be$, accounting for slow radioactive decays. All optional weak rate corrections in PRIMAT (see Ref. Pitrou:2018cgg) and the new corrections from neutrino spectral distortions are included in the calculations, see text for more details. The $68\%$ constraints on the helium abundance $Y_{P}$ from the CMB ($Y_{P}=0.242\pm 0.012$) and the recent EMPRESS experiment ($Y_{P}=0.2370^{+0.0034}_{-0.0033}$) are indicated by the light and dark blue shaded regions, respectively.
@@
Figure 5 illustrates the resulting light element abundances, incorporating all weak rate corrections from (80) and the QED corrections up to $\mathcal{O}(e^{3})$ for plasma thermodynamics and certain nuclear rates, as a function of the neutrino degeneracy parameter $\xi_{\nu}$. The abundance of ${}^{3}{\rm He}$ refers to ${}^{3}{\rm He}+\rm T$, while ${}^{7}{\rm Li}$ corresponds to ${}^{7}{\rm Li}+\rm^{7}Be$, accounting for slow radioactive decays. Since the abundances of different elements can vary by several orders of magnitude, we normalize the element abundances to their respective values at $\xi_{\nu}=0$. Consequently, all curves in Fig. 5 intersect at $\xi_{\nu}=0$. We observe that the abundances of $\rm D,\,^{3}{\rm He},\,^{4}{\rm He}$, and ${}^{7}{\rm Li}$ decrease with increasing $\xi_{\nu}$ due to the corresponding reduction in the neutron-to-proton number density ratio. Furthermore, the variation in ${}^{4}{\rm He}$ abundance is most sensitive to changes in $\xi_{\nu}$, while the ${}^{3}{\rm He}$ abundance is the least sensitive. Additionally, Fig. 5 shows the constraints on helium abundance derived from the CMB and EMPRESS experiments:
$Y_{P}|_{\rm CMB}=0.242\pm 0.012,\quad Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033},$ (83)
@@
as represented by the light and dark blue shaded regions. Our results indicate that the constraints on $Y_{P}$ within the $68\%$ confidence interval can be translated into constraints on the neutrino degeneracy parameter $\xi_{\nu}$:
${\rm CMB}:\quad-0.026\leq$ $\,\xi_{\nu}\leq 0.069\,,$ (84)
${\rm EMPRESS}:\quad 0.027\leq$ $\,\xi_{\nu}\leq 0.054\,.$ (85)
@@
Clearly, the EMPRESS experiment favors a non-zero positive $\xi_{\nu}$, while the CMB experiment remains compatible with the Standard Model case of $\xi_{\nu}=0$.
@@
Figure 6: The relative difference in light element abundances compared to the results in the neutrino instantaneous decoupling limit as a function of the neutrino degeneracy parameter $\xi_{\nu}$. The solid and dotted curves correspond to results with and without weak corrections induced by neutrino spectral distortions, respectively. The color scheme for the curves is consistent with that in Fig. 5, except that negative results for ${}^{7}{\rm Li}$ are presented here for simplicity.
@@
Table 1: Light element abundances, including all weak rate corrections in (80), for various implementations of neutrino decoupling and degeneracy parameters. ${}^{3}{\rm He}$ denotes $(^{3}{\rm He}+\mathrm{T})$ and ${}^{7}{\rm Li}$ denotes $(^{7}{\rm Li}+\rm Be)$ to account for slow radioactive decays.
@@
Table 2: Relative differences in light element abundances, comparing cases with and without corrections induced by neutrino spectral distortions to results in the neutrino instantaneous decoupling limit. ${}^{3}{\rm He}$ denotes $(^{3}{\rm He}+\mathrm{T})$ and ${}^{7}{\rm Li}$ denotes $(^{7}{\rm Li}+\rm Be)$ to account for slow radioactive decays.
@@
It is also interesting to discuss the effects of corrections induced by neutrino spectral distortions on BBN. We examine the impact of different treatments of the neutrino decoupling process, considering three implementations: (1) the neutrino instantaneous decoupling limit, $i.e.,$ , $N_{\rm eff}^{\rm inst}$, as discussed in the previous section, where thermal Fermi-Dirac distributions for neutrinos and antineutrinos are used in the BBN calculations; (2) the no-SD case, which employs the actual $N_{\rm eff}$ and neutrino temperatures, incorporating the non-instantaneous decoupling effect while only using thermal Fermi-Dirac distributions for neutrinos and antineutrinos. This treatment is the most common in current BBN studies; (3) the actual SD case, which utilizes the actual $N_{\rm eff}$, accounting for the non-instantaneous decoupling effect and incorporating actual density distributions with spectral distortions for neutrinos and antineutrinos. This case corresponds to the results shown in Fig. 5.
Using the BBN results in the instantaneous decoupling limit as a benchmark, we present the relative differences for the no-SD and actual-SD cases, indicated by dotted and solid curves in Fig. 6, respectively. Note that we have plotted the negative of the relative differences for ${}^{7}{\rm Li}$ in Fig. 6 due to the specific nuclear reaction rates utilized in PRIMAT. We report the explicit abundances of helium-4, deuterium, helium-3, and lithium-7 for the degeneracy parameters $\xi_{\nu}=0,\pm 1$ in table 1, along with the associated relative variations compared to the baseline abundance in table 2. Our results with $\xi_{\nu}=0$ are consistent with those reported in Froustey:2020mcq , which also considered the effects of neutrino spectral distortions on BBN within the standard model (SM). As shown in Fig. 6 and the accompanying tables, the actual-SD and no-SD cases yield nearly identical light element abundances for $\xi_{\nu}=-0.1$. The BBN abundances in the actual-SD case show minimal dependence on $\xi_{\nu}$, while the traditional no-SD treatment predicts a stronger dependence for both elements. Notably, the no-SD case predicts negative variations for the helium abundance with positive $\xi_{\nu}$ compared to the instantaneous decoupling limit. In contrast, the more precise results with spectral distortions indicate positive variations in helium abundance, demonstrating that incorporating neutrino spectral distortions is crucial for accurate predictions of BBN abundances.
@@
Finally, we provide fitting functions for the light element abundances as follows:
@@
where the subscript $\xi_{\nu}=0$ indicates results obtained in the SM case using actual density matrices, as detailed in Table 1. These fitting functions facilitate further studies, yielding fractional errors smaller than $0.01\%$ within the range $-0.1\leq\xi_{\nu}\leq 0.1$.  [^3] The coefficients of our fitting functions differ slightly from those in Pitrou:2018cgg ; Froustey:2024mgf , which is mainly due to the different neutron lifetimes we have chosen. In addition, our fitting function is more precise.
5 Impacts of Primordial Neutrino Asymmetry on the CMB and LSS
In this section, we examine the effects of primordial neutrino asymmetries on the evolution of the late universe, particularly focusing on the cosmic microwave background (CMB) and large-scale structure (LSS). Primordial neutrino asymmetry affects the CMB and LSS in three primary ways:
@@ 5 Impacts of Primordial Neutrino Asymmetry on the CMB and LSS
    The expansion rate of the universe is altered by the increased $N_{\rm eff}$ resulting from primordial neutrino asymmetries. Specifically, the additional radiation density delays the time of matter-radiation equality, enhancing the early integrated Sachs-Wolfe effect in the CMB angular power spectrum.
@@
    As shown in Fig. 5, the electron neutrino asymmetry significantly modifies the helium abundance, which subsequently affects the late-time evolution of the universe. This impacts features such as the tail of the CMB angular power spectrum due to diffusion damping and the baryon acoustic oscillations (BAO).
@@
    Additionally, to account for neutrino oscillations during neutrino decoupling, neutrino masses must be considered. These masses influence the late-time evolution of the universe, particularly affecting matter-radiation equality, even in the standard model (SM) scenario with thermal Fermi-Dirac distributions. For the SM case, CMB and BAO observations can place stringent constraints on the sum of neutrino masses, yielding $\sum m_{\nu}<0.07\,{\rm eV}$ DESI:2024mwx . We will revisit these constraints in the next section. Furthermore, the neutrino Boltzmann hierarchy for massive neutrinos is modified by the neutrino degeneracy parameter and spectral distortions.
In the following, we first review the impact of massive neutrinos on the universe and present the Boltzmann hierarchies for electrons, photons, and neutrinos. We then discuss the effects of primordial neutrino asymmetries on CMB temperature anisotropies and the linear matter power spectrum, computed using the Einstein-Boltzmann code CLASS Lesgourgues:2011re ; Lesgourgues:2011rh .
5.1 Effects of Primordial Neutrino Asymmetries in Structure Formation
To accurately describe flavor oscillations during neutrino decoupling, neutrino mass must be considered, which has significant implications for structure formation. Massive neutrinos are the only species in the SM that behave as radiation in the early universe and as matter in the late universe.
In the early universe, neutrinos and antineutrinos remain relativistic after decoupling, as long as their effective temperatures are much higher than their masses. In this regime, their energy densities contribute to the total radiation energy density, parameterized by $N_{\rm eff}$. As the universe expands and the effective temperatures decrease as $a^{-1}$, neutrinos and antineutrinos become non-relativistic when their effective temperatures fall below their masses, at which point they contribute to the matter content. The transition to the non-relativistic phase and the free-streaming effects of massive neutrinos leave noticeable imprints on CMB anisotropies and LSS.
As discussed, primordial neutrino asymmetries modify this scenario. In addition to indirect changes from the altered Hubble rate and helium abundance, we emphasize several direct effects of primordial neutrino asymmetries.
@@ 5.1.1 Effects on Neutrino Abundance
First, the total relic abundance of neutrinos and antineutrinos increases directly due to primordial neutrino asymmetries. At present, the total neutrino relic abundance is given by
@@
where $y=p/T_{\rm CM}$ is the comoving momentum defined in section 3, $T_{\rm CM,0}$ is the current comoving temperature, $\rho_{\rm crit}$ is the critical energy density, and $\varrho_{i}(y)$ and $\bar{\varrho}_{i}(y)$ represent the comoving density distributions of neutrinos and antineutrinos, respectively, for $i=1,2,3$, corresponding to the three mass eigenstates with mass $m_{i}$.
The present neutrino abundance is further modified by spectral distortions of neutrinos and antineutrinos, though the integrals over their distributions cannot be carried out analytically in this case. The effect of primordial neutrino asymmetries becomes clearer when spectral distortions are neglected, allowing an analytical evaluation of the integrals, as shown in the second row of (90). The first factor, $\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}$, represents the usual neutrino relic abundance in the SM, while the second factor accounts for the effects of neutrino asymmetries, with ${\rm Li}_{j}(x)$ being the polylogarithms of order $j$ and $\zeta(x)$ the Riemann zeta function. In the SM case, ${\rm Li}_{3}(-1)=3\zeta(3)/4$, so the second factor equals 1. For non-zero $\xi_{\nu}$, the second factor depends on the absolute value of $\xi_{\nu}$, with larger $\xi_{\nu}$ leading to a larger factor and, consequently, a larger neutrino abundance. However, the corrections from primordial neutrino asymmetries are small. For instance, the relative difference in $\Omega_{\nu}h^{2}$ for $\xi_{\nu}=0.2$ is only $1.54\%$. Similarly, the corrections to the neutrino abundance from spectral distortions of neutrinos and antineutrinos are negligible.